Mr. Roberts bought a gas cooker for $945. He sold it to a costumer for $803.25 due to a damage. [Loss or profit
Answer by math_tutor2020(3835)


cost = 945
revenue = 803.25

profit = revenue - cost
profit = 803.25 - 945
profit = -141.75
The negative profit indicates a loss

Loss percentage = (lossAmount/cost)*100%
Loss percentage = (141.75/945)*100%
Loss percentage = 0.15*100%
Loss percentage = 15%
The cooker lost 15% of its value when Mr Roberts sold it.

Or note how
revenue/cost = 803.25/945 = 0.85
Meaning that he was able to sell it for 85% of its original cost, therefore it lost 100% - 85% = 15% of its value.